OpenStudy (cookiemonster18):

1.The table below shows the height (in inches) and weight (in pounds)of eight basketball players. Height=67, 69, 70, 72, 74, 78, 79 Weight=183, 201, 206, 240, 253, 255 what is the correlation of the set of data? Round your answer to the nearest thousandth. A.-0.946 B.0.596 C.0.035 D.0.981

OpenStudy (solomonzelman):

you only have 6 weights

OpenStudy (solomonzelman):

please check your question

OpenStudy (cookiemonster18):

sorry, here the correct table to the question @SolomonZelman Height=67, 69, 70, 72, 74, 74, 78, 79 Weight=183, 201, 206, 220, 226, 240, 253, 255

OpenStudy (cookiemonster18):

@Cardinal_Carlo can you help with this question?

OpenStudy (anonymous):

Correlation is Positive when the values increase together, and Negative when one value decreases as the other increases. In our case, weight increases as height increases. So, our correlation value is positive. We can cross out A.

OpenStudy (anonymous):

Correlation is zero (or at least, near zero) when the values exhibit no trend. However, in our case, this is not true. Again, weight increases as height increases. So, our correlation is not anywhere near zero. We can cross C out.

OpenStudy (cookiemonster18):

so we cross out A we're left with B and D

OpenStudy (anonymous):

That's right. First, let's draw an average line through our two endpoints. \[m = \frac{ 255-183 }{ 79-67 } = \frac{ 72 }{ 12 } = 6\] \[183 = 6(67) + b\] \[b = -219\] And, we would get this average line: \[w = f(h) = 6h - 219\] To test if there is a high or low correlation (alternatively, we're saying, 'is it D or B?' respectively), you have to compare your points to this average line. If the values stay within proximity of the average line, then there's a high correlation. Contrarily, if the points stray too far from this average line, then there's a low correlation. TIP: It's best if you mark these points on a representative graph @CookieMonster18

OpenStudy (cookiemonster18):

im i correct that it's D. 0.981

OpenStudy (anonymous):

You have a good chance that you might be. This question is actually referring to an algebraic approach to determine correlation. However, it is long and complicated to explain fully in detail. But, if you'd like to take a look here are its steps: Step 1: Find the mean of x, and the mean of y Step 2: Subtract the mean of x from every x value (call them "a"), do the same for y (call them "b") Step 3: Calculate: a × b, a2 and b2 for every value Step 4: Sum up a × b, sum up a2 and sum up b2 Step 5: Divide the sum of a × b by the square root of [(sum of a2) × (sum of b2)] If you prefer looking at its formula form, it's attached below.

OpenStudy (anonymous):

Your professor gave you a busy work question. My sympathies.